# Learn With Imagination: Safe Set Guided State-wise Constrained Policy Optimization

### v3

@@ 1 Introduction
Meanwhile, safe control methods to continuously meet stringent safety requirements in predictable environments are widely examined, with energy function-based methods being the most prevalent Khatib (1986); Ames et al. (2014); Liu & Tomizuka (2014); Gracia et al. (2013); Wei & Liu (2019). These methods establish energy functions assigning lower energy to safe states and project nominal control into energy dissipating control, hence persistently maintaining system safety. However, these approahces heavily rely on accessible white-box analytical models of system dynamics. The Implicit Safe Set Algorithm (ISSA) Zhao et al. (2021) addresses this problem using black-box optimization over black-box dynamics models. Nevertheless, the reliance on continuous monitoring to ensure safety imposes significant computational overhead, and the RL policy itself does not inherently guarantee safe behavior. This combination presents a major obstacle to the successful deployment of the algorithm in real-time, safety-critical applications.
@@ 3.1 Safe Reinforcement Learning
Existing safe RL approaches either consider safety after convergence or safety during training Zhao et al. (2023c). End-to-end approaches are usually used to ensure safety after convergence Liang et al. (2018); Tessler et al. (2018); Bohez et al. (2019); Ma et al. (2021); He et al. (2023b). However, these approaches cannot avoid unsafe explorations.
Safety in training is achieved by hierarchical approaches which uses a safeguard to filter out unsafe explorative actions. The safeguard relies on the knowledge on the system dynamics, which can be either learned dynamics Dalal et al. (2018); Thananjeyan et al. (2021); Zhao et al. (2022), white-box dynamics Fisac et al. (2018); Shao et al. (2021), or black-box dynamics Zhao et al. (2021). Nevertheless, these methods cannot guarantee convergence of the RL policy.
@@ 3.2 Implicit Safe Set Algorithm
Energy function-based methods  Wei & Liu (2019) achieve safe control by designing an energy function offline such that 1) the low energy states are safe and 2) there always exists a feasible control input to dissipate the energy. A typical design Liu & Tomizuka (2014) was proposed as $\phi=\phi_{0}^{*}+k_{1}\dot{\phi}_{0}+\cdots+k_{n}\phi_{0}^{(n)}$. It is shown in Liu & Tomizuka (2014) that if the control input is unbounded ($\mathcal{A}=\mathbb{R}^{n_{a}}$), then there always exist a control input that satisfies the constraint $\dot{\phi}\leq 0\text{ when }\phi=0$; and if the control input always satisfies that constraint, then the set $\bar{\mathcal{S}}:=\{s\mid\phi(s)\leq 0\}\cap\{s\mid\phi_{0}(s)\leq 0\}$ is forward invariant. In practice, the actual control signal is computed through a quadratic projection of the nominal control $a^{r}$ to the control constraint
$a=$ $\mathop{\mathbf{argmin}}_{a\in\mathcal{A}}\|a-a^{r}\|^{2}\text{ s.t. }\dot{\phi}\leq-\eta(\phi),$ (5)
where $\eta(\phi)$ is designed to be a positive constant when $\phi\geq 0$ and $\infty$- ∞ when $\phi<0$ according to SSA Liu & Tomizuka (2014). Here, we define the set of safe control as $\mathcal{A}_{S}(s):=\{a\in\mathcal{A}\mid\dot{\phi}\leq-\eta(\phi)\}$. To match the notion of MDP, we should consider discrete-time safe control set $\mathcal{A}_{S}^{D}(s):=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq\max\{\phi(s)-\eta,0\}\}$.
Based on these, the implicit safe set algorithm (ISSA) Zhao et al. (2021) was proposed to construct the black-box dynamics based safeguard ensuring persistent satisfication of safety specification. The black-box dynamics can be a digital twin or neural network. Under some mild assumptions, it first synthesizes a safety index to make sure $\mathcal{A}_{S}(s)$ is nonempty for all $s$ (details of the mild assumptions and safety index synthesis are summarized in Appendix A). Then it manages to project the reference action generated from RL policy $\pi_{\theta}$ into $\mathcal{A}_{S}(s)$ during policy training. In detail, the nominal control $a_{t}^{r}$ needs to be projected to $\mathcal{A}_{S}^{D}(s_{t})$ by solving the following optimization:
@@
To solve (6), they first introduce a sample-efficient Adaptive Momentum Boundary Approximation (AdamBA) algorithm. Then, ISSA directly uses it to find the safe control with minimum deviation from the reference control along the boundaries of $\mathcal{A}_{S}^{D}(s)$. If the process fails to return a solution, grid sampling will be deployed to find a safe control; and AdamBA is deployed again to improve the solution optimality with respect to (6). We summarize details of AdamBA and ISSA in Algorithm 2 and Algorithm 3, respectively.
@@ 4.2 S-3PO
where $\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]$ is KL divergence between two policy $(\pi^{\prime},\pi)$ at state $\hat{s}$, the set $\{\pi\in\Pi_{\theta}~{}:~{}\mathbb{E}_{\hat{s}\sim\bar{d}^{\pi_{k}}}[\mathcal{D}_{KL}(\pi\|\pi_{k})[\hat{s}]]\leq\delta\}$ is called trust region, $d^{\pi_{k}}\doteq(1-\gamma)\sum_{t=0}^{H}\gamma^{t}P(\hat{s}_{t}=\hat{s}|{\pi_{k}})$, $\bar{d}^{\pi_{k}}\doteq\sum_{t=0}^{H}P(\hat{s}_{t}=\hat{s}|{\pi_{k}})$ and $\epsilon^{{\pi}}_{D}\doteq\mathbf{max}_{\hat{s}}|\mathbb{E}_{a\sim{\pi}}[A^{\pi_{k}}_{D}(\hat{s},a)]|$.
Remark 3.
@@
5 Practical Implementation
Weighted loss for cost value targets
Figure 2: $V_{[D_{i}]\pi}(\hat{s})$ target of five sampled episodes.
Line search scheduling
@@ 6 Theoretical Results
In this section, We present three theorems, including (i) zero violation exploration (Theorem 1), (ii) state-wise safety guarantee of S-3PO (Theorem 2), and (iii) worst case performance guarantee of S-3PO (Theorem 3). The proofs of the three theorems are summarized in Appendix C, Appendix D and Appendix E, respectively.
Theorem 1 (Zero Training Time Violation).
@@
Theorem 2 (Safety Guarantee of S-3PO).
@@
Theorem 3 (Worst-case Performance Degradation in S-3PO.).
@@
Remark 4.
Theorem 1 shows zero safety violation exploration by proving the system state will never leave $\mathcal{\bar{S}}$ under the safeguard of ISSA. Theorem 2 shows that by satisfying the constraint of (11), the new policy is guaranteed to generate safe action, i.e. $a\in\mathcal{A}_{S}^{D}$, in expectation. Intuitively, Theorem 3 shows that with enough training epochs, the reward performance of S-3PO will not deteriorate too much after each update.
@@ 7 Experiments
In our experiments, we aim to answer the following questions:
Q1: Does S-3PO achieve zero-violation during the training?
Q2: How does S-3PO without safeguard compare with other state-of-the-art safe RL methods?
Q3: Does S-3PO learn to act without safeguard?
Q4: How does weighted loss trick impact the performance of S-3PO?
Q5: Is “imaginary” cost necessary to achieve zero violation?
Q6: How does S-3PO scale to high dimensional robots?
@@ 7.1 Experiments Setup
(a) Point
(b) Swimmer
(c) Ant
(d) Hazard
(e) Pillar
Figure 3: Robots and constraints for benchmark problems in our environment.
Environment Setting
@@
Our experiments revolve around four distinct robotic entities that can be categorized into two primary types:
  1. 1.
    Wheel Robot: this category encompasses robots that use wheels for locomotion, maintaining a continuous and seamless interaction with their surrounding environment. An example is the Point in fig. 3(a) which is designated as $\mathcal{A}\subseteq\mathbb{R}^{2}$.
  2. 2.
    Link Robot: this group includes robots composed of multiple connected links, which interact intermittently with their surroundings through the extremities of these links. Furthermore, the shape of these robots can undergo dynamic changes during these interactions. Specifically, in our environment, we have 1) Swimmer shown in fig. 3(b) as a three-link robot ($\mathcal{A}\subseteq\mathbb{R}^{2}$); 2) Ant shown in fig. 3(c) as a quadrupedal robot ($\mathcal{A}\subseteq\mathbb{R}^{8}$).
Two different types of constraints are considered.
  1. 1.
    Hazard: Dangerous areas shown in fig. 3(d). Hazards are trespassable circles on the ground. The agent is penalized for entering them.
  2. 2.
    Pillar: Fixed obstacles shown in  fig. 3(e). The agent is penalized for hitting them.
    More details about the experiments are discussed in Section F.1.
(a) Point_1Hazard
(b) Point_8Hazard
(c) Point_1Pillar
(d) Swimmer_1Hazard
Figure 5: Zero violation performance from four representative test suites in low dimensional systems (evaluated with the safeguard).
Comparison Group
The methods in the comparison group include: (i) unconstrained RL algorithm TRPO (Schulman et al., 2015) and TRPO-ISSA. (ii) end-to-end constrained safe RL algorithms CPO (Achiam et al., 2017), TRPO-Lagrangian (Bohez et al., 2019), PCPO (Yang et al., 2020), SCPO (Zhao et al., 2023b). (iii) We select TRPO as our baseline method since it is state-of-the-art and already has safety-constrained derivatives that can be tested off-the-shelf. For all experiments, the policy $\pi$, the value $(V^{\pi},V_{D}^{\pi})$ are all encoded in feedforward neural networks using two hidden layers of size (64,64) with tanh activations. The full list of parameters of all methods compared can be found in Section F.2.
Evaluation Metrics
For comparison, we evaluate algorithm performance based on (i) reward performance, (ii) average episode cost and (iii) cost rate. More details are provided in Section F.3. We set the limit of cost to 0 for all the safe RL algorithms since we aim to avoid any violation of the constraints.
(a) Point_1Hazard
(b) Point_8Hazard
(c) Point_1Pillar
(d) Swimmer_1Hazard
Figure 7: Results from four representative test suites in low dimensional systems (evaluated without the safeguard).
(a) Point_1Hazard
(b) Point_8Hazard
(c) Point_1Pillar
(d) Swimmer_1Hazard
Figure 9: Triggering frequency of the safeguard from four representative test suites in low dimensional systems.
@@ 7.2 Evaluating S-3PO and Comparison Analysis
The training performance of four representative test suites are summarized in Figure 5, where the S-3PO algorithm clearly outperforms other baseline methods by achieving zero violations, consistent with the safety guarantee outlined in Theorem 1. For more experiments, please check Appendix F. This superior performance is attributed to the safeguard mechanism within the S-3PO framework, which effectively corrects unsafe actions at every step, particularly during training. Furthermore, as demonstrated in Figure 7, the reward performance continues to improve throughout the learning process, remaining comparable to state-of-the-art baselines. This distinct capability of S-3PO ensures safe reinforcement learning with zero safety violations, even in real-world scenarios, addressing Q1.
State-wise Safety Without Safety Monitor
@@
To assess the performance of the S-3PO policy without the safeguard, evaluations are conducted at the end of each epoch. During these evaluations, the S-3PO policy is tested over 10,000 steps without the safeguard. This allows us to determine whether S-3PO effectively learns a state-wise safe policy through the guidance of the safe set-guided cost. The comparative results are presented in Figure 7. When contrasted with other baseline safe RL approaches, S-3PO demonstrates superior performance even without the safeguard, achieving (i) near-zero average episode cost and (ii) significantly reduced cost rates, all while maintaining competitive reward performance. These findings are consistent with Theorem 2 and Theorem 3, highlighting that by minimizing imaginary safety violations, the policy rapidly learns to act safely, which addresses Q2.
Learn to Act without Safeguard
@@
As highlighted in 1, the key concept behind penalizing imaginary safety violations is to minimize the activation of the safeguard, thereby significantly reducing its computational complexity and enabling real-time implementation. To illustrate this, we visualize the average number of times the ISSA-based safeguard is triggered per step in Figure 9. For comparison, TRPO-ISSA is included as a baseline, which relies continuously on the safeguard to maintain safe control. Figure 9 shows that S-3PO dramatically reduces the frequency of safeguard activations, approaching zero, indicating that a state-wise safe policy has been effectively learned, thus addressing Q3.
Ablation on Weighted Loss for Fitting Cost Increment Value Targets
@@
As pointed in Section 5, fitting $V_{D_{i}}(\hat{s}_{t})$ is a critical step towards solving S-3PO, which is challenging due to non-increasing stair shape of the target sequence. To elucidate the necessity of weighted loss for solving this challenge, we evaluate the cost rate of S-3PO under six distinct weight settings (0.0, 0.2, 0.4, 0.6, 0.8, 1.0) on Point_4Hazard test suite. The results shown in Figure 11(a) validates that a larger weight (hence higher penalty on predictions that violate the characteristics of value targets) results in better cost rate performance. This ablation study answers Q4.
(a) Comparison of cost rate performance with 6 different weights.
(b) Comparison between “imaginary” cost and action correction cost.
Figure 11: Comparison results of S-3PO
Necessity of “Imaginary” Cost
@@
To understand the importance of the "imaginary" cost within the S-3PO framework, we compare it to another cost based on the magnitude of action correction Chen & Liu (2021). This empirical analysis is conducted using the Point_4Hazard test suite. As shown in Figure 11(b), the "imaginary" cost yields superior cost rate performance. This suggests that the "imaginary" cost offers deeper insights into the complex dynamics between the robot and its environment, thereby addressing Q5.
Figure 13: Cost and cost rate performance of Ant_1Hazard.
Scale S-3PO to High-Dimensional Link Robots
@@
To showcase S-3PO’s scalability and performance with complex, high-dimensional link robots, we conducted additional tests on Ant_1Hazard featuring 8 dimensional control spaces. As shown in Figure 13, S-3PO effectively drives the cost to zero and rapidly reduces the cost rate, showcasing its clear superiority in high-dimensional safety policy learning and highlighting its exceptional scalability to more complex systems.

### v4

@@ 1 Introduction
Meanwhile, safe control methods to continuously meet stringent safety requirements in predictable environments are widely examined, with energy function-based methods being the most prevalent Khatib (1986); Ames et al. (2014); Liu and Tomizuka (2014); Gracia et al. (2013); Wei and Liu (2019). These methods establish energy functions assigning lower energy to safe states and project nominal control into energy dissipating control, hence persistently maintaining system safety. However, these approahces heavily rely on accessible white-box analytical models of system dynamics. The Implicit Safe Set Algorithm (ISSA) Zhao et al. (2021, 2024) addresses this problem using black-box optimization over black-box dynamics models. Nevertheless, the reliance on continuous monitoring to ensure safety imposes significant computational overhead, and the RL policy itself does not inherently guarantee safe behavior. This combination presents a major obstacle to the successful deployment of the algorithm in real-time, safety-critical applications.
@@ 3.1 Safe Reinforcement Learning
Existing safe RL approaches either consider safety after convergence or safety during training Zhao et al. (2023c). End-to-end approaches are usually used to ensure safety after convergence Liang et al. (2018); Tessler et al. (2018); Bohez et al. (2019); Ma et al. (2021); He et al. (2023b). However, these approaches cannot avoid unsafe explorations. Safety in training is achieved by hierarchical approaches which uses a safeguard to filter out unsafe explorative actions. The safeguard relies on the knowledge on the system dynamics, which can be either learned dynamics Dalal et al. (2018); Thananjeyan et al. (2021); Zhao et al. (2022), white-box dynamics Fisac et al. (2018); Shao et al. (2021), or black-box dynamics Zhao et al. (2021). Nevertheless, these methods cannot guarantee convergence of the RL policy.
@@ 3.2 Implicit Safe Set Algorithm
The Implicit Safe Set Algorithm (ISSA) Zhao et al. (2021, 2024) ensures the persistent satisfaction of safety specifications for systems with black-box dynamics (e.g., digital twins or neural networks) through energy function-based optimization. Leveraging energy function $\phi=\phi_{0}^{*}+k_{1}\dot{\phi}_{0}+\cdots+k_{n}\phi_{0}^{(n)}$ and theoretical results from SSA Liu and Tomizuka (2014), ISSA synthesizes a safety index to guarantee that the safe control set $\mathcal{A}_{S}(s):=\{a\in\mathcal{A}\mid\dot{\phi}\leq-\eta(\phi)\}$ is nonempty. And then the set $\bar{\mathcal{S}}:=\{s\mid\phi(s)\leq 0\}\cap\{s\mid\phi_{0}(s)\leq 0\}$ is forward invariant under mild assumptions. Here $\eta(\phi)$ is designed to be a positive constant when $\phi\geq 0$ and $\infty$- ∞ when $\phi<0$. Intuitively, the energy function $\phi$ ensures that low-energy states are safe, while the control signal within $\mathcal{A}_{S}(s)$ reduces $\phi$ whenever the system state is outside the safe region. This further derives the forward invariance of $\bar{\mathcal{S}}$, meaning that once the system enters the safe region, its state will stays within the safe part of the state space.
To match the notion of MDP, we consider discrete-time safe control set $\mathcal{A}_{S}^{D}(s):=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq\max\{\phi(s)-\eta,0\}\}$. Thus ISSA ensures safety by projecting the nominal control $a_{t}^{r}$ generated from RL policy $\pi_{\theta}$ to $\mathcal{A}_{S}^{D}(s_{t})$ through solving the following optimization:
@@
The detailed mild assumoptions and safety index synthesis are summarized in Appendix A. And the algorithms for solving (5) are provided in Algorithms 2 and 3.
@@ 4.2 S-3PO
where $\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]$ is KL divergence between two policy $(\pi^{\prime},\pi)$ at state $\hat{s}$, the set $\{\pi\in\Pi_{\theta}~{}:\\ ~{}\mathbb{E}_{\hat{s}\sim\bar{d}^{\pi_{k}}}[\mathcal{D}_{KL}(\pi\|\pi_{k})[\hat{s}]]\leq\delta\}$ is called trust region, $d^{\pi_{k}}\doteq(1-\gamma)\sum_{t=0}^{H}\gamma^{t}P(\hat{s}_{t}=\hat{s}|{\pi_{k}})$, $\bar{d}^{\pi_{k}}\doteq\sum_{t=0}^{H}P(\hat{s}_{t}=\hat{s}|{\pi_{k}})$ and $\epsilon^{{\pi}}_{D}\doteq\mathbf{max}_{\hat{s}}|\mathbb{E}_{a\sim{\pi}}[A^{\pi_{k}}_{D}(\hat{s},a)]|$.
Remark 4.3.
@@
4.3 Practical Implementation
The pseudocode of S-3PO is give as algorithm 1. Here we summarize two techniques that helps with S-3PO’s practical performance. (i) Weighted loss for cost value targets: A critical step in S-3PO involves fitting the cost increment value function, $V_{D}^{\pi}(\hat{s}_{t})$, which represents the maximum future cost increment relative to the highest state-wise cost observed so far. This function follows a non-increasing staircase pattern along the trajectory. Thus, we adopt a weighted loss function, $L_{weight}$, to penalize predictions that violate the non-increasing property: $L_{weight}=L(\hat{y}_{t}-y_{t})*(1+w*\text{1}[(\hat{y}_{t}-y_{t-1})>0])$, where $L$ denotes Mean Squared Error, $\hat{y}_{t}$ is the prediction, ${y}_{t}$ is the fitting target and ${w}$ is the penalty weight. (ii) Line Search scheduling: Constraints in (10) might become infeasible due to approximation errors. In this case, we perform a recovery update, enforcing the cost advantage $A_{D}^{\pi}$ to decrease from early training steps $k_{\mathrm{safe}}$ while focusing on reward improvements of $A^{\pi}$ towards the end of training, prioritizing safety first and reward performance later. Please check Section F.3 for furthur details.
@@ 5 Theoretical Results
In this section, We present three theorems, including (i) zero violation exploration (Theorem 5.1), (ii) state-wise safety guarantee of S-3PO (Theorem 5.2), and (iii) worst case performance guarantee of S-3PO (Theorem 5.3). The proofs of the three theorems are summarized in Appendix C, Appendix D and Appendix E, respectively.
Theorem 5.1 (Zero Training Time Violation).
@@
Theorem 5.2 (Safety Guarantee of S-3PO).
@@
Theorem 5.3 (Worst-case Performance Degradation in S-3PO.).
@@
Remark 5.4.
Theorem 5.1 shows zero safety violation exploration by proving the system state will never leave $\mathcal{\bar{S}}$ under the safeguard of ISSA. Theorem 5.2 shows that by satisfying the constraint of (10), the new policy is guaranteed to generate safe action, i.e. $a\in\mathcal{A}_{S}^{D}$, in expectation. Intuitively, Theorem 5.3 shows that with enough training epochs, the reward performance of S-3PO will not deteriorate too much after each update.
@@ 6 Experiments
In our experiments, we aim to answer: Q1: Does S-3PO achieve zero-violation during the training? Q2: How does S-3PO without safeguard compare with other state-of-the-art safe RL methods? Q3: Does S-3PO learn to act without safeguard? Q4: How does weighted loss trick impact the performance of S-3PO? Q5: Is “imaginary” cost necessary to achieve zero violation? Q6: How does S-3PO scale to high dimensional robots?
@@ 6.1 Experiments Setup
Comparison Group The methods in the comparison group include: (i) unconstrained RL algorithm TRPO (Schulman et al., 2015) and TRPO-ISSA. (ii) end-to-end constrained safe RL algorithms CPO (Achiam et al., 2017), TRPO-Lagrangian (Bohez et al., 2019), PCPO (Yang et al., 2020), SCPO (Zhao et al., 2023b). (iii) We select TRPO as our baseline method since it is state-of-the-art and already has safety-constrained derivatives that can be tested off-the-shelf. The full list of parameters of all methods compared can be found in Section F.2.
Metrics For comparison, we evaluate algorithm performance based on (i) reward performance, (ii) average episode cost and (iii) cost rate. More details are provided in Section F.4. We set the limit of cost to 0 for all the safe RL algorithms since we aim to avoid any violation of the constraints.
@@ 6.2 Evaluating S-3PO and Comparison Analysis
Zero Violation During Training The training performance of four representative test suites are summarized in LABEL:fig:zero_violation_performance, where the S-3PO algorithm clearly outperforms other baseline methods by achieving zero violations, consistent with the safety guarantee outlined in Theorem 5.1. For more experiments, please check Appendix F. This superior performance is attributed to the safeguard mechanism within the S-3PO framework, which effectively corrects unsafe actions at every step, particularly during training. Furthermore, as demonstrated in LABEL:fig:comparison_results-main, the reward performance remains comparable to state-of-the-art baselines. This distinct capability of S-3PO ensures safe reinforcement learning with zero safety violations, addressing Q1.
@@
State-wise Safety Without Safety Monitor At the end of each epoch, the S-3PO policy is tested over 10,000 steps without the safeguard. This allows us to determine whether S-3PO effectively learns a state-wise safe policy through the guidance of the safe set-guided cost. As shown in in LABEL:fig:comparison_results-main, S-3PO demonstrates superior performance even without the safeguard, achieving (i) near-zero average episode cost and (ii) significantly reduced cost rates, all while maintaining competitive reward performance. These findings are consistent with Theorem 5.2 and Theorem 5.3, highlighting that by minimizing imaginary safety violations, the policy rapidly learns to act safely, which addresses Q2.
@@
Learn to Act without Safeguard As highlighted in 1, the key concept behind penalizing imaginary safety violations is to minimize the activation of the safeguard, thereby significantly reducing its computational complexity and enabling real-time implementation. To illustrate this, we visualize the average number of times the ISSA-based safeguard is triggered per step in LABEL:fig:comparison_results_ISSA-main. For comparison, TRPO-ISSA is included as a baseline, which relies continuously on the safeguard to maintain safe control. LABEL:fig:comparison_results_ISSA-main shows that S-3PO dramatically reduces the frequency of safeguard activations, approaching zero, indicating that a state-wise safe policy has been effectively learned, thus addressing Q3.
@@
Ablation on Weighted Loss for Fitting Cost Increment Value Targets As pointed in Section 4.3, fitting $V_{D_{i}}(\hat{s}_{t})$ is a critical step towards solving S-3PO, which is challenging due to non-increasing stair shape of the target sequence. To elucidate the necessity of weighted loss for solving this challenge, we evaluate the cost rate of S-3PO under six distinct weight settings (0.0, 0.2, 0.4, 0.6, 0.8, 1.0) on Point_4Hazard test suite. The results shown in LABEL:fig:weight_performance validates that a larger weight (hence higher penalty on predictions that violate the characteristics of value targets) results in better cost rate performance. This ablation study answers Q4.
@@
Necessity of “Imaginary” Cost To understand the importance of the "imaginary" cost within the S-3PO framework, we compare it to another cost based on the magnitude of action correction Chen and Liu (2021). This empirical analysis is conducted using the Point_4Hazard test suite. As shown in LABEL:fig:action_correction_performance, the "imaginary" cost yields superior cost rate performance. This suggests that the "imaginary" cost offers deeper insights into the complex dynamics between the robot and its environment, thereby addressing Q5.
Figure 5: Cost performance of Ant_1Hazard.
@@
Scale S-3PO to High-Dimensional Link Robots To showcase S-3PO’s scalability and performance with complex, high-dimensional link robots, we conducted additional tests on Ant_1Hazard featuring 8 dimensional control spaces. As shown in Figure 5, S-3PO effectively drives the cost to zero and rapidly reduces the cost rate, showcasing its superiority in high-dimensional safety policy learning and highlighting its exceptional scalability to more complex systems, which answers Q6.